# Boosted GFlowNets Improving Exploration via Sequential Learning

### v1

@@ A Monte Carlo estimator with $K$ reverse paths is
Under the notations of LABEL:eq:interp-tb, let $Z_{\mathrm{old}}$ and $Z_{\theta}$ be the normalizing constants for $\hat{R}_{\mathrm{old}}(x)\coloneqq\mathbb{E}[\hat{R}_{\mathrm{old}}(x;\tau^{\prime})]$ and $\hat{R}_{\theta}(x;\tau)$. Also, let $\beta=\nicefrac{{Z_{\theta}}}{{Z_{\mathrm{old}}+Z_{\theta}}}$ be the mixture weight in Equation˜15. Then, define $p_{\mathrm{old}}(\tau)\propto\hat{R}_{\mathrm{old}}(x)p_{B}(\tau|x)$ and $p_{\mathrm{tgt}}(\tau)=R(x)p_{B}(\tau|x)$ as the trajectory-level distributions induced by our current and target models, and let $p_{M}(\tau)=(1-\beta)\cdot p_{\mathrm{old}}(\tau)+\beta\cdot p_{F}(\tau)$ be the corresponding mixture distribution. In this scenario,
@@
$L_{\mathrm{boost}}(\tau)=\left(\log\frac{\hat{R}(x;\theta)+\hat{R}_{\mathrm{old}}(x)}{R(x)}\right)^{2}.$ (30)
@@
$L_{\mathrm{boost}}(\tau)=\left(\log\frac{Z_{\theta}p_{F}(\tau)+\hat{R}_{\mathrm{old}}(x)p_{B}(\tau|x)}{R(x)p_{B}(\tau|x)}\right)^{2}.$ (31)
By definition, $Z_{\mathrm{old}}$ is the normalizing constant for $\hat{R}_{\mathrm{old}}(x)$. Then, under the notations of our proposition,
$L_{\mathrm{boost}}(\tau)=\left(\log\frac{\beta p_{F}(\tau)+(1-\beta)p_{\mathrm{old}}(\tau)}{\frac{R(x)}{Z_{\mathrm{old}}+Z_{\theta}}p_{B}(\tau|x)}\right)^{2}=\left(\log\frac{\beta p_{F}(\tau)+(1-\beta)p_{\mathrm{old}}(\tau)}{p_{\mathrm{tgt}}(\tau)}\right)^{2}.$ (32)
@@
$\nabla_{\theta}\mathcal{D}_{\mathrm{KL}}[p_{M}||p_{\mathrm{tgt}}]=\underset{{\tau\sim p_{M}}}{\mathbb{E}}\left[\nabla_{\theta}\log p_{M}(\tau)+\left(\log\frac{p_{M}(\tau)}{p_{\mathrm{tgt}}(\tau)}\right)\cdot\nabla_{\theta}\log p_{M}(\tau)\right]=\underset{{\tau\sim p_{M}}}{\mathbb{E}}\left[\left(\log\frac{p_{M}(\tau)}{p_{\mathrm{tgt}}(\tau)}\right)\cdot\nabla_{\theta}\log p_{M}(\tau)\right],$ (35)
in which we used the fact that $\mathbb{E}_{\tau\sim p}[\log p(\tau)]=0$ for any probability distribution $p$ (williams1992).

### v2

@@ Appendix D Theorems and Proofs
Let $\widehat{Z}$ and $Z_{\theta}$ be the normalizing constants for $\widehat{R}(x)\coloneqq\mathbb{E}_{\tau\sim P_{B}(\cdot\mid x)}[\widehat{Z}\frac{\widehat{P}_{F}(\tau)}{P_{B}(\tau\mid x)}]$ and ${R}_{\theta}(x;\tau)\coloneq Z_{\theta}\frac{P_{F}(\tau)}{P_{B}(\tau\mid x)}$. Also, let $\beta=\nicefrac{{Z_{\theta}}}{{\widehat{Z}+Z_{\theta}}}$ be the mixture weight in Equation˜15. Then, define $\widehat{p}(\tau)\propto\widehat{R}(x)p_{B}(\tau|x)$ and $p_{\mathrm{tgt}}(\tau)=R(x)p_{B}(\tau|x)$ as the trajectory-level distributions induced by our current and target models, and let $p_{M}(\tau)=(1-\beta)\cdot\widehat{p}(\tau)+\beta\cdot p_{F}(\tau)$ be the corresponding mixture distribution. In this scenario,
@@
$L_{\mathrm{boost}}(\tau)=\left(\log\frac{{R}(x;\theta)+\widehat{R}(x)}{R(x)}\right)^{2}.$ (34)
@@
$L_{\mathrm{boost}}(\tau)=\left(\log\frac{Z_{\theta}p_{F}(\tau)+\widehat{R}(x)p_{B}(\tau|x)}{R(x)p_{B}(\tau|x)}\right)^{2}.$ (35)
By definition, $\widehat{Z}$ is the normalizing constant for $\widehat{R}(x)$. Then, under the notations of our proposition,
$L_{\mathrm{boost}}(\tau)=\left(\log\frac{\beta p_{F}(\tau)+(1-\beta)\widehat{p}(\tau)}{\frac{R(x)}{\widehat{Z}+Z_{\theta}}p_{B}(\tau|x)}\right)^{2}=\left(\log\frac{\beta p_{F}(\tau)+(1-\beta)\widehat{p}(\tau)}{p_{\mathrm{tgt}}(\tau)}\right)^{2}.$ (36)
@@
$\nabla_{\theta}\mathcal{D}_{\mathrm{KL}}[p_{M}||p_{\mathrm{tgt}}]$ $=\underset{{\tau\sim p_{M}}}{\mathbb{E}}\left[\nabla_{\theta}\log p_{M}(\tau)+\left(\log\frac{p_{M}(\tau)}{p_{\mathrm{tgt}}(\tau)}\right)\cdot\nabla_{\theta}\log p_{M}(\tau)\right]$ (39)
$=\underset{{\tau\sim p_{M}}}{\mathbb{E}}\left[\left(\log\frac{p_{M}(\tau)}{p_{\mathrm{tgt}}(\tau)}\right)\cdot\nabla_{\theta}\log p_{M}(\tau)\right],$
in which we used the fact that $\mathbb{E}_{\tau\sim p}[\log p(\tau)]=0$ for any probability distribution $p$ (Williams, 1992).